6. Consider the following supply and demand curves for bat teries: where PB is the cost of batteries and Pc is the cost of copper. a) Determine the equilibrium price and quantity for bat- teries (PR and QR) in terms of the variable Poc b) Using comparative statics, discuss how the equilib- rium Pß and QB changes as Pc increases and de creases c) Determine the price elasticity of supply and price elas- ticity of demand as a function of PB a d) Let Pe 2, for what values of PB is the price elastic- itv of demand elastic, unitarv elastic, and inelastic! e) Determine the cross-price elasticity of demand for bat- teries with respect to the price for copper. є08P what can we say about copper and batteries (demand substitutes or complements)?
Answer #1

a) The supply and demand equations are given as follows:

QS = 10 +8PB +7PC ------(i) &

QD =100-4PB -2PC ..........(ii)

Under market equilibrium

QS = QD

or, 10 +8PB +7PC = 100-4PB -2PC

or 12PB +9PC =90

or, 4PB = 30-3PC

or PB = 3/4(10-PC )

Putting the value of PB in equation (i) or (ii) we get,

QB= QS =QD = 10+8.3/4(10-PC ) + 7PC = 70+PC

b) From part a)

equilibrium price (PB ) = 3/4(10-PC )

& equilibrium quantity (QB ) = 70 +PC

Here when value of PC increases, the equilibrium price (PB ) decreases and equilibrium quantity (QB ) increases.

But when value of PC decreases, the equilibrium price (PB ) increases and equilibrium quantity (QB ) decreases.

c) Price elasticity of demand (eD ) = dQD /dPB * PB /QD

                                             = -4*PB / (100-4PB -2PC )

                                         = 4PB /(4PB +2PC -100)

Price elasticity of supply (es ) = dQS /dPB *PB /QS

                                         = 8PB /(10+8PB +7PC )

d) If PC = 2 then demand and supply equations is

QD = 100-4PB -2*2= 96-4PB &

QS = 10 +8PB +7*2= 24+8PB

Therefore, elasticity of demand (eD ) = dQD /dPC * PB /QD = -2PC /(96-4PB ) = -4/(96-4PB )

for unitary elastic demand, eD = 1

or, -4/(96-4PB ) = 1 or, 4PB = 100 and PB =25

For inelastic demand, eD = 0 or, 96-4PB =0 or, PB =24

For elastic demand, eD >0

Putting PB =26 we get

eD = -4/(96-4*26) = 4/8=1/2>0

e) Cross elasticity of demand for batteries with respect to price for copper is

eCD = dQD /dPC *PC /QD = -2PC /(100-4PB -2PC )